OpenStudy (goformit100):

If the baseball team wins 78 of the 114 ga mes that it plays, what percent has it lost?

OpenStudy (goformit100):

@Koikkara

OpenStudy (koikkara):

arithmetic....

OpenStudy (koikkara):

hmm.....got the trick....from thepdf... http://www.lenoircc.edu/pdf/LCC_BasicArith_review.pdf

OpenStudy (anonymous):

total games=114 games won=78 games lost=114-78=36 % games lost=(36*100)/114=600/19